A project requires an investment of $15000 today and it is expected to generate free cash flows of $5000 per year for 4 years. The company’s weighted average cost of capital is 10.1% per year. What is the projects equivalent annual annuity?
Answer #1

NPV = ΣCFn/(1+r)n, where CFn is the cash flow for year n

r = cost of capital = 10.1%

CF0 = -15000
CF1 = 5000
CF2 = 5000
CF3 = 5000
CF4 = 5000

=> NPV = -15000 + 5000/(1+0.101) + 5000/(1+0.101)2 + 5000/(1+0.101)3 + 5000/(1+0.101)4

= $815.08

Let the equivalent annual annuity be P

=> NPV = P/(1+0.101) + P/(1+0.101)2 + P/(1+0.101)3 + P/(1+0.101)4 = P[1- (1+0.101)-4]/0.101

=> P[1- (1+0.101)-4]/0.101 = 815.08

=> P = $257.69